# MML Probabilistic Principal Component Analysis

### v3

@@ I Introduction
  • We show that the MML threshold for detecting a latent factor agrees with the Baik-Ben Arous-Péché (BBP) phase transition threshold [25]. The MML threshold is slightly higher than the theoretical distinguishability limit to prevent false positives caused by finite-sample fluctuations of the residual variance estimate.
@@ III Minimum message length analysis of the PCA model
$r({\bf x})=\sum_{\theta\in\Theta}\pi(\theta)\,p({\bf x}|\theta).$ (14)
@@ III-C Prior information
$\pi(\bm{\alpha},\sigma,\bm{\phi})=\pi_{\sigma}(\sigma)\,\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})|J_{{\bf A}\to\bm{\phi}}|\,J!,$ (41)
@@ III-D Codelength
where $A=N(K-J)-KJ,B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. The solution space of the MML estimate exhibits a phase-transition behavior that can be classified into two regimes:
@@
$\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}=\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$ (55)
@@
The derivation uses first-order perturbation theory around the maximum likelihood root. All expectations are taken under fixed $K$, $J$ and $N\to\infty$, and higher-order covariance terms between the sample eigenvalues and $\hat{\tau}_{\rm ML}$ are $O(N^{-2})$ and therefore neglected. The MML estimate of $\tau$ is a stationary point of the polynomial
@@
The coefficients of this polynomial converge such that a root is exactly $\hat{\tau}_{ML}$ for $N\to\infty$. For finite $N$, we will approximate the MML estimate of $\tau$ as the MLE estimate plus a small correction term $\Delta$. The bias of the maximum likelihood estimate of $\tau$ up to second order [44] is
$\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}=-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+O(N^{-2}).$ (56)
Let $A(\tau)$ denote the characteristic polynomial of the $J$ sample eigenvalues
@@
$Q(\tau)$ $=\sum_{k=0}^{J}(-1)^{k+2}\left(\frac{k(K-J+1)-KJ}{N(K-J)}\right)e_{J-k}\tau^{k+1}$
$=\frac{\tau}{N(K-J)}\sum_{k=0}^{J}(-1)^{k}\left(k(K-J+1)-KJ\right)e_{J-k}\tau^{k}$
$=\frac{\tau}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau)-KJA(\tau)\right).$
@@
$\Delta$ $=-\frac{\frac{\tau_{\rm ML}}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau_{\rm ML})-KJA(\tau_{\rm ML})\right)}{A(\hat{\tau}_{\rm ML})}$
$=\frac{\hat{\tau}_{\rm ML}}{N(K-J)}\left(KJ-(K-J+1)\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}\right).$
Next, replace sample quantities with their population counterparts and compute the expectation
$\mathbb{E}\{\Delta\}$ $=\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right).$
Recall that the MML estimate of $\tau$ is modelled as the MLE plus a small correction. We finally combine expectations and simplify to obtain the bias estimate up to second order:
$\mathbb{E}\{\hat{\tau}_{\rm MML}\}$ $=\mathbb{E}\{\hat{\tau}_{\rm MLE}\}+\mathbb{E}\{\Delta\}$
$=\tau-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2})$
$=\tau+\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$
The maximum likelihood estimate is negatively biased and underestimates noise, with bias of order $O(1/N)$ and approximately proportional to $J\tau/N$. In contrast, the MML estimate of $\tau$ is positively biased and overestimates noise in finite samples, with bias of order $O(1/(NK))$ and approximately proportional to $J^{2}\tau/(NK)$. The absolute ratio $R$ of the two biases is:
$R=\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-J)\left(\frac{J+\sum_{j=1}^{J}\rho_{j}^{-1}}{J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right).$ (60)
Observe that the MML estimate reduces bias compared to the maximum likelihood estimate by a factor roughly proportional to the dimension $K$. If the signals are strong (high signal-to-noise ratio) with $\rho_{j}\to\infty$ ($j=1,\ldots,J)$, the ratio is approximately $R\approx(K-J)/J$, suggesting that the MML estimate reduces bias by a factor proportional to the ratio of the total dimension to the latent dimension. In contrast, for weak signals with $\rho_{j}\to 0$, the terms $\rho_{j}^{-1}$ dominate and the bias reduction factor for the MML estimate is $R\approx(K-J)$. Thus, even in the case of weak signals, where the bias of both MML and maximum likelihood estimate is inflated, the relative advantage of the MML estimate remains constant. Next, we analyse model selection properties of the MML codelength.
@@
$\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-1)\left(\frac{1+\sum_{j=1}^{J}\rho_{j}^{-1}}{1^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right)=K-1.$ (68)
For one true latent factor, the MML estimate reduces bias by the factor of (K1)(K-1) irrespective of the signal-to-noise ratio. When the signal is very weak, both biases blow up in magnitude because the spike is barely separated from the noise bulk. The maximum likelihood estimate severely underestimates $\tau$, while MML slightly overestimates it, but by a factor 1/(K1)1/(K-1). When the signal is very strong, both biases shrink to $O(1/N)$ constants with the maximum likelihood bias being negative and about (K1)(K-1) times larger than the MML bias.

### v4

@@ I Introduction
  • We show that the MML threshold for detecting a latent factor agrees with the Baik-Ben Arous-Péché (BBP) phase transition threshold [25].
MATLAB code implementing the methodology discussed in this paper that can be used to reproduce all experiments in Section IV is available in our github repository.
@@ III Minimum message length analysis of the PCA model
$r({\bf x})=\int_{\theta\in\Theta}\pi(\theta)\,p({\bf x}|\theta)d\theta.$ (14)
@@ III-C Prior information
$\pi(\bm{\alpha},\sigma,\bm{\phi})=\pi_{\sigma}(\sigma)\,\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})\pi_{H}({\bf H}_{B})|J_{{\bf A}\to\bm{\phi}}|\,J!,$ (41)
@@
Remark. In the classical regime ($N\to\infty$, $K$ fixed), finite‑sample PCA eigenvectors concentrate around the population eigenvectors at a rate governed by $N$ and the spike signal-to-noise ratio $\rho_{j}$ $(j\leq J)$. Following the perturbation analysis in [44], the principal angle $\theta_{j}$ between the sample PCA eigenvector and the true (population) eigenvector satisfies
$\mathbb{E}\left\{\sin^{2}\theta_{j}\right\}\asymp\frac{K-1}{N\rho_{j}},\quad\rho_{j}=\alpha_{j}^{2}/\sigma^{2}.$
This may be incorporated this into the orientation prior $\pi_{H}(\cdot)$ by replacing the uniform Stiefel prior with a matrix Langevin prior conditional on an unknown mean orientation $H_{0}$, with concentration $\kappa_{j}=N\rho_{j}$:
$\pi_{H}({\bf H}_{B}\mid\alpha,\tau)=c_{K,J}(\bm{\kappa})\,\exp\!\Big\{\mathrm{tr}\!\big(\bm{\kappa}\,{\bf H}_{0}^{\prime}{\bf H}_{B}\big)\Big\},\qquad{\bf H}_{0}\sim\mathrm{Unif}\!\left(V_{J}(\mathbb{R}^{K})\right),$ (42)
where $H_{0}$ is the unobserved mean orientation with a uniform prior and $c_{K,J}(\bm{\kappa})$ is a normalising constant. This approach preserves rotational invariance of the prior while encoding the finite‑sample resolution of the PCA directions.
@@ III-D Codelength
where $A=N(K-J)-KJ,B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. The solution space of the MML estimate of $\tau$ exhibits a phase-transition behavior that can be classified into two regimes:
@@
$\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}=\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+\frac{\tau}{N^{2}(K-J)}\mathcal{B}(K,J,\{\rho_{j}\})+O(N^{-3})$ (56)
where $\mathcal{B}:=\mathcal{B}(K,J,\{\rho_{j}\})$ is
$\mathcal{B}=-\Big(J+\sum_{m=1}^{J}\rho_{m}^{-1}\Big)\Big[KJ+2(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-2}\Big]-(K-J+1)\sum_{j=1}^{J}\frac{(b_{j}/\tau)}{\rho_{j}^{2}},$ (57)
and $b_{j}$ is the $O(N^{-1})$ finite-sample mean-shift coefficient of the $j$-th signal eigenvalue, given explicitly by
$\frac{b_{j}}{\tau}=(1+\rho_{j})\left[\sum_{\ell\leq J,\ \ell\neq j}\frac{1+\rho_{\ell}}{\rho_{j}-\rho_{\ell}}+\frac{K-J}{\rho_{j}}\right].$ (58)
@@
The derivation uses first-order perturbation theory around the maximum likelihood root. All expectations are taken under fixed $K$, $J$ and $N\to\infty$ From [45] and [44], we have for finite NN
$\delta_{j}=\lambda_{j}+\frac{b_{j}}{N}+O(N^{-2}),\quad j\leq J,$
where $\lambda_{j}$ is the corresponding population signal eigenvalue and $b_{j}$ $(j\leq J)$ is the $O(N^{-1})$ bias coefficient defined in (58). The expected value of the sample signal eigenvalue $\delta_{j}$ [44] is
$\mathbb{E}\left\{\delta_{j}\right\}$ $=\lambda_{j}+\frac{1}{N}\tau(1+\rho_{j})\left[\sum_{\ell\leq J,\ \ell\neq j}\frac{1+\rho_{\ell}}{\rho_{j}-\rho_{\ell}}+\frac{K-J}{\rho_{j}}\right]=\lambda_{j}+\frac{b_{j}}{N}+O(N^{-2}),\quad j\leq J.$
Subsequently, the maximum likelihood estimate of $\tau$ can be written as
$\hat{\tau}_{\rm ML}=\tau+\frac{u}{N}+O(N^{-2}),\quad u=-\tau\sum_{j=1}^{J}(1+\rho_{j}^{-1}),$ (59)
and the the finite-sample bias of the MLE of $\tau$ (see also [46]) is
$\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}=\frac{u}{N}+O(N^{-2})=-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+O(N^{-2}),$ (60)
up to second order. Recall that the MML estimate of $\tau$ is a stationary point of the polynomial
@@
The coefficients of this polynomial converge such that a root is exactly $\hat{\tau}_{ML}$ for $N\to\infty$. For finite $N$, we will approximate the MML estimate of $\tau$ as the MLE estimate plus a small correction term $\Delta$. Let $A(\tau)$ denote the characteristic polynomial of the $J$ sample eigenvalues
@@
$Q(\tau)$ $=\sum_{k=0}^{J}(-1)^{k+2}\left(\frac{k(K-J+1)-KJ}{N(K-J)}\right)e_{J-k}\tau^{k+1}=\frac{\tau}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau)-KJA(\tau)\right).$
@@
$\Delta$ $=\frac{\hat{\tau}_{\rm ML}}{N(K-J)}\left(KJ-(K-J+1)\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}\right),$
up to $O(N^{-2})$ remainder terms. A first‑order Taylor expansion in $1/N$ gives
$T=\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}$ $=-\sum_{j=1}^{J}\rho_{j}^{-1}+\frac{1}{N\tau}\sum_{j=1}^{J}\left[-\frac{u}{\rho_{j}}-\frac{u-b_{j}}{\rho_{j}^{2}}\right]+O(N^{-2}),$
where $u=-\tau\sum_{m=1}^{J}(1+\rho_{m}^{-1})$, as before. Substituting (59) and $T$ into $\Delta$ and simplifying, we get
$\mathbb{E}\left\{\Delta\right\}$ $=\frac{\tau}{N(K-J)}\left[KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right]+\frac{\tau}{N^{2}(K-J)}\mathcal{B}(K,J,\{\rho_{j}\})+O(N^{-3}),$
where $\mathcal{B}$ is given in (57). Recall that the MML estimate of $\tau$ is modelled as the MLE plus a small correction
$\mathbb{E}\{\hat{\tau}_{\rm MML}\}$ $=\mathbb{E}\{\hat{\tau}_{\rm MLE}\}+\mathbb{E}\{\Delta\}.$
Substituting $\mathbb{E}\left\{\Delta\right\}$ and (60) into the above expression and simplifying completes the proof. ∎
Up to second order, the maximum likelihood estimate is negatively biased and underestimates noise, with bias of order $O(1/N)$ approximately proportional to $J\tau/N$. In contrast, the MML estimate of $\tau$ is positively biased and overestimates noise in finite samples, with bias of order $O(1/(NK))$ approximately proportional to $J^{2}\tau/(NK)$. The absolute ratio $R$ of the two biases is:
$R_{N}=\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-J)\left(\frac{J+\sum_{j=1}^{J}\rho_{j}^{-1}}{(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1})+\mathcal{B}/N}\right).$ (64)
Observe that the MML estimate reduces bias compared to the maximum likelihood estimate by a factor roughly proportional to the dimension $K$. If the signals are strong (high signal-to-noise ratio) with $\rho_{j}\to\infty$ ($j=1,\ldots,J)$, the ratio $R_{N}$ is
$R_{N}=\frac{K-J}{J}\left(1+\frac{K}{N}+O(N^{-2})\right),$ (65)
which is approximately $R\approx(K-J)/J$. This suggests that the MML estimate reduces bias by a factor proportional to the ratio of the total dimension to the latent dimension. In contrast, for weak signals with $\rho_{j}\to 0$, the terms $\rho_{j}^{-1}$ dominate and the bias reduction factor for the MML estimate is $R\approx(K-J)$. However, in this regime the $J$-factor MML stationary point typically ceases to exist and the MML procedure rejects the $J$-factor model in favor of J1J-1. Next, we analyse model selection properties of the MML codelength.
@@
$\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-1)\frac{1}{1-\frac{K(K+\rho_{1}(\rho_{1}+2))}{N\rho_{1}^{2}}}+O(N^{-2}),$ (73)
where $\rho_{1}=\alpha_{1}^{2}/\sigma^{2}$ is the signal-to-noise ratio for the top factor. As $\rho_{1}\to\infty$, the MML estimate of $\tau$ reduces bias by (K1)/(1K/N)(K-1)/(1-K/N) (to second order). When the signal is very weak, both biases blow up in magnitude because the spike is barely separated from the noise bulk. The maximum likelihood estimate severely underestimates $\tau$, while MML slightly overestimates it, but by a factor 1/(K1)1/(K-1). When the signal is very strong, both biases shrink to $O(1/N)$ constants with the maximum likelihood bias being negative and about (K1)(K-1) times larger than the MML bias.
@@ V Discussion
VI Data Availability
The data in this article are available in the MML-PCA GitHub repository at: https://github.com/EnesMakalic/MML-PCA.